You have already invested $300 in a stock with an annual return of 11%. How much of an additional $1,300 should be invested at 12% and how much at 6% so that the total return on the entire $1,600 is 9%? (Round each answer to the nearest cent.)
Answer by JoelSchwartz(130)

x=amount invested at twelve percent
y=is equal to the amount at six percent.
I know that the amount at twelve percent plus the amount at six percent is equal to 1300.
x+y=1300
We also know that three hundred dollars is invested plus thirteen hundred dollars which equals 1600 dollars. The interest rate times sixteen hundred is the total interest of the investment.
.09*1600=144
I know that three hundred times eleven percent is equal to 33
.11*300=33
33+.12x+.06y=144
.12x=144-33-.06y
.12x=111-.06y
x=925-1/2y
x+y=1300
925-1/2y+y=1300
925+1/2y=1300
1/2y=375
y=750
x=925-1/2*750
x=925-375
x=550